Tara and Ben each run 6km every morning. Tara runs at an average speed of 10 km/h, while Ben runs at an average speed of 3 m/s.How long after Tara, should Ben begin running if they wish to complete their runs at the same time? Give your answer in minutes and seconds.
Answer by ankor@dixie-net.com(22740)

Tara and Ben each run 6km every morning.
Tara runs at an average speed of 10 km/h, while Ben runs at an average speed of 3 m/s.
How long after Tara, should Ben begin running if they wish to complete their runs at the same time?
Give your answer in minutes and seconds.
:
Find how long it takes T to run 6 km
t = 6/10 hr
t .6(60) = 36 min or 36*60 = 2160 seconds
:
Find how long it takes B to run 6 km
6 km = 6000 meters
t = 6000/3
t = 2000 seconds
:
2160 - 2000 = 160 seconds, which is 2 min 40 sec
:
Ben should begin 2 min 40 sec after Tara starts